Chemical cold packs contain two pouches, one containing solid ammonium nitrate and the other containing water (often dyed for no real reason). When the contents of the pouches mix, the ammonium nitrate dissolves in the water, NH4NO3(s) → NH4+(aq) + NO3-(aq)
The dissolving process is endothermic. The reaction absorbs approximately 326 J/g of ammonium nitrate. To lower the temperature of 220. g of water by 20.0°C, how many grams of ammonium nitrate would you need to dissolve?
1
Expert's answer
2015-11-02T12:24:11-0500
To determine the necessary amount of amounium nitrate, we can use the heat balance of water:
Q = cmΔT and Q = 326x, where x is a mass of ammonium nitrate.
The final equation is cmΔT = 326x;
x = cmΔT/326 = 4200*0.22*20/326 = 56.69 g

Answer: 56.69 g